OpenStudy (anonymous):

Out of 100 students sampled, 70 of them said that they hoped to get married someday. With 99.7% confidence, what is the approximate percentage of the students in the population who hope to get married someday? A. 51.7% to 88.3% B. 60.8% to 79.2% C. 65.4% to 74.6% D. 56.3% to 83.7%

OpenStudy (anonymous):

the sample proportion is \(\hat p=70/100=0.7\) and the corresponding sample proportion distribution's standard deviation is \(\sigma_{\hat p}=\sqrt{\frac1N\hat p(1-\hat p)}\approx0.0458\). 99.7% confidence intervals cover values up to 3 standard deviations from the mean by the empirical rule, so our interval should be approximately \(\hat p\pm3\sigma_{\hat p}\approx0.7\pm0.1375=(0.5625,0.8375)\) i.e. the interval is between \(56.3\%\) and \(83.7\%\)

OpenStudy (anonymous):

thank you